You are given the system of equations

"\begin{array}{ccc}2x & +4y & =10\\ 3x & +2y & =7 \end{array} ."

Which of the following is true about the system above?


  1. The determinant of the coefficient matrix is -8
  2. With Cramer's rule, the value of x is "x=-\\frac{\\left| \\begin{array}{cc}10 & 4\\\\ 7 & 2 \\end{array} \\right |}{8}.\u200b"
  3. With Cramer's rule, the value of y is "y=-\frac{\left| \begin{array}{cc}2 & 10\\ 3 & 7 \end{array} \right |}{8}."
  4. The determinant of the coefficient matrix is 8
  5. With Cramer's rule, the value of y is "y=\frac{\left| \begin{array}{cc}2 & 10\\ 3 & 7 \end{array} \right |}{8}."
1
Expert's answer
2021-09-16T07:53:02-0400

since the determinant of the coefficient matrix is "\left| {\begin{matrix}
2&4\\
3&2
\end{matrix}} \right| = 4 - 12 = - 8" then statement 1 is true. Then then statement 2 and 3 are true and statement 4 and 5 are false.

Answer: 1, 2, 3 are true, 4 and 5 are false